\begin{thebibliography}{10}
\expandafter\ifx\csname url\endcsname\relax
  \def\url#1{\texttt{#1}}\fi
\expandafter\ifx\csname urlprefix\endcsname\relax\def\urlprefix{URL }\fi
\expandafter\ifx\csname href\endcsname\relax
  \def\href#1#2{#2} \def\path#1{#1}\fi

\bibitem{mccurdy1984}
A.~McCurdy, K.~C. Ng, B.~N. Parlett,
  \href{https://doi.org/10.2307/2008291}{Accurate computation of divided
  differences of the exponential function}, Math. Comput. 43~(168) (1984)
  501--528.
\newblock \href {https://doi.org/10.2307/2008291} {\path{doi:10.2307/2008291}}.
\newline\urlprefix\url{https://doi.org/10.2307/2008291}

\bibitem{deboor2005}
C.~de~Boor,
  \href{https://surveys-in-approximation-theory.com/papers/2/index.html}{Divided
  {{Differences}}}, Surv. Approximation Theory 1 (2005) 46--69.
\newline\urlprefix\url{https://surveys-in-approximation-theory.com/papers/2/index.html}

\bibitem{DaleckiiKrein}
Y.~L. Daleckii, M.~G. Krein, Stability of Solutions of Differential Equations
  in Banach Space, Vol.~43 of Translations of Mathematical Monographs, American
  Mathematical Society, Providence, RI, 1974.

\bibitem{BirmanSolomyak1968}
M.~S. Birman, M.~Z. Solomyak, Double stieltjes operator integrals, in:
  M.~Shimada (Ed.), Topics in Mathematical Physics, Vol.~2, Consultants Bureau,
  New York, 1968, pp. 19--46, english translation of the 1966 Leningrad Univ.
  paper.
\newblock \href {https://doi.org/10.1007/978-1-4684-7592-0_3}
  {\path{doi:10.1007/978-1-4684-7592-0_3}}.

\bibitem{BhatiaMatrixAnalysis}
R.~Bhatia, Matrix Analysis, Vol. 169 of Graduate Texts in Mathematics,
  Springer, New York, 1997.
\newblock \href {https://doi.org/10.1007/978-1-4612-0653-8}
  {\path{doi:10.1007/978-1-4612-0653-8}}.

\bibitem{Chernoff1952}
H.~Chernoff, A measure of asymptotic efficiency for tests of a hypothesis based
  on the sum of observations, Annals of Mathematical Statistics 23~(4) (1952)
  493--507.
\newblock \href {https://doi.org/10.1214/aoms/1177729330}
  {\path{doi:10.1214/aoms/1177729330}}.

\bibitem{Hoeffding1963}
W.~Hoeffding, Probability inequalities for sums of bounded random variables,
  Journal of the American Statistical Association 58~(301) (1963) 13--30.
\newblock \href {https://doi.org/10.1080/01621459.1963.10500830}
  {\path{doi:10.1080/01621459.1963.10500830}}.

\bibitem{Tropp2012}
J.~A. Tropp, User-friendly tail bounds for sums of random matrices, Foundations
  of Computational Mathematics 12~(4) (2012) 389--434.
\newblock \href {https://doi.org/10.1007/s10208-011-9099-z}
  {\path{doi:10.1007/s10208-011-9099-z}}.

\bibitem{Parlett1998}
B.~N. Parlett, The Symmetric Eigenvalue Problem, Society for Industrial and
  Applied Mathematics, Philadelphia, PA, 1998.
\newblock \href {https://doi.org/10.1137/1.9781611971163}
  {\path{doi:10.1137/1.9781611971163}}.

\bibitem{Higham2008}
N.~J. Higham, Functions of Matrices: Theory and Computation, Society for
  Industrial and Applied Mathematics, Philadelphia, PA, 2008.
\newblock \href {https://doi.org/10.1137/1.9780898717778}
  {\path{doi:10.1137/1.9780898717778}}.

\bibitem{VanLoan1978}
C.~F. Van~Loan, Computing integrals involving the matrix exponential, IEEE
  Transactions on Automatic Control 23~(3) (1978) 395--404.
\newblock \href {https://doi.org/10.1109/TAC.1978.1101743}
  {\path{doi:10.1109/TAC.1978.1101743}}.

\bibitem{CoxMatthews2002}
S.~M. Cox, P.~C. Matthews, Exponential time differencing for stiff systems,
  Journal of Computational Physics 176~(2) (2002) 430--455.
\newblock \href {https://doi.org/10.1006/jcph.2002.6995}
  {\path{doi:10.1006/jcph.2002.6995}}.

\bibitem{KassamTrefethen2005}
A.-K. Kassam, L.~N. Trefethen, Fourth-order time-stepping for stiff pdes, SIAM
  Journal on Scientific Computing 26~(4) (2005) 1214--1233.
\newblock \href {https://doi.org/10.1137/S1064827502410633}
  {\path{doi:10.1137/S1064827502410633}}.

\bibitem{HochbruckOstermann2010}
M.~Hochbruck, A.~Ostermann, Exponential integrators, Acta Numerica 19 (2010)
  209--286.
\newblock \href {https://doi.org/10.1017/S0962492910000048}
  {\path{doi:10.1017/S0962492910000048}}.

\bibitem{NiesenWright2012}
J.~Niesen, W.~M. Wright, Algorithm 919: A krylov subspace algorithm for
  evaluating the phi-functions appearing in exponential integrators, ACM
  Transactions on Mathematical Software 38~(3) (2012) 1--19.
\newblock \href {https://doi.org/10.1145/2168773.2168781}
  {\path{doi:10.1145/2168773.2168781}}.

\bibitem{CaliariKandolfOstermannRainer2016LejaBEA}
M.~Caliari, P.~Kandolf, A.~Ostermann, S.~Rainer, The leja method revisited:
  Backward error analysis for the matrix exponential, SIAM Journal on
  Scientific Computing 38~(3) (2016) A1639--A1661.
\newblock \href {https://doi.org/10.1137/15M1027620}
  {\path{doi:10.1137/15M1027620}}.

\bibitem{KandolfOstermannRainer2014LejaResidual}
P.~Kandolf, A.~Ostermann, S.~Rainer, A residual based error estimate for leja
  interpolation of matrix functions, Linear Algebra and its Applications 458
  (2014) 157--173.
\newblock \href {https://doi.org/10.1016/j.laa.2014.04.023}
  {\path{doi:10.1016/j.laa.2014.04.023}}.

\bibitem{Guettel2013RationalKrylovReview}
S.~G{\"u}ttel, Rational krylov approximation of matrix functions: Numerical
  methods and optimal pole selection, GAMM-Mitteilungen 36~(1) (2013) 8--31.
\newblock \href {https://doi.org/10.1002/gamm.201310002}
  {\path{doi:10.1002/gamm.201310002}}.

\bibitem{Peller2006MOI}
V.~V. Peller, Multiple operator integrals and higher operator derivatives,
  Journal of Functional Analysis 233~(2) (2006) 515--544.
\newblock \href {https://doi.org/10.1016/j.jfa.2005.09.003}
  {\path{doi:10.1016/j.jfa.2005.09.003}}.

\bibitem{Peller2016MOISurvey}
V.~V. Peller, Multiple operator integrals in perturbation theory, Bulletin of
  Mathematical Sciences 6~(1) (2016) 15--88.
\newblock \href {https://doi.org/10.1007/s13373-015-0073-y}
  {\path{doi:10.1007/s13373-015-0073-y}}.

\bibitem{barash2020}
L.~Gupta, L.~Barash, I.~Hen,
  \href{https://linkinghub.elsevier.com/retrieve/pii/S0010465520301673}{Calculating
  the divided differences of the exponential function by addition and removal
  of inputs}, Comput. Phys. Commun. 254 (2020) 107385.
\newblock \href {https://doi.org/10.1016/j.cpc.2020.107385}
  {\path{doi:10.1016/j.cpc.2020.107385}}.
\newline\urlprefix\url{https://linkinghub.elsevier.com/retrieve/pii/S0010465520301673}

\bibitem{Caliari2007DividedDifferences}
M.~Caliari, Accurate evaluation of divided differences for polynomial
  interpolation of exponential propagators, Computing 80~(2) (2007) 189--201.
\newblock \href {https://doi.org/10.1007/s00607-007-0227-1}
  {\path{doi:10.1007/s00607-007-0227-1}}.

\bibitem{Zivcovich2019}
F.~Zivcovich, \href{http://drna.padovauniversitypress.it/2019/1/4}{Fast and
  accurate computation of divided differences for analytic functions, with an
  application to the exponential function}, Dolomites Research Notes on
  Approximation 12~(1) (2019) 28--42.
\newblock \href {https://doi.org/10.14658/pupj-drna-2019-1-4}
  {\path{doi:10.14658/pupj-drna-2019-1-4}}.
\newline\urlprefix\url{http://drna.padovauniversitypress.it/2019/1/4}

\bibitem{albash2017}
T.~Albash, G.~Wagenbreth, I.~Hen,
  \href{https://link.aps.org/doi/10.1103/PhysRevE.96.063309}{Off-diagonal
  expansion quantum {{Monte Carlo}}}, Phys. Rev. E 96~(6) (2017) 063309.
\newblock \href {https://doi.org/10.1103/PhysRevE.96.063309}
  {\path{doi:10.1103/PhysRevE.96.063309}}.
\newline\urlprefix\url{https://link.aps.org/doi/10.1103/PhysRevE.96.063309}

\bibitem{barash2024}
L.~Barash, A.~Babakhani, I.~Hen,
  \href{https://link.aps.org/doi/10.1103/PhysRevResearch.6.013281}{Quantum
  {{Monte Carlo}} algorithm for arbitrary spin-1/2 {{Hamiltonians}}}, Phys.
  Rev. Res. 6~(1) (2024) 013281.
\newblock \href {https://doi.org/10.1103/PhysRevResearch.6.013281}
  {\path{doi:10.1103/PhysRevResearch.6.013281}}.
\newline\urlprefix\url{https://link.aps.org/doi/10.1103/PhysRevResearch.6.013281}

\bibitem{gupta2020}
L.~Gupta, T.~Albash, I.~Hen,
  \href{https://iopscience.iop.org/article/10.1088/1742-5468/ab9e64}{Permutation
  matrix representation quantum {{Monte Carlo}}}, J. Stat. Mech. 2020~(7)
  (2020) 073105.
\newblock \href {https://doi.org/10.1088/1742-5468/ab9e64}
  {\path{doi:10.1088/1742-5468/ab9e64}}.
\newline\urlprefix\url{https://iopscience.iop.org/article/10.1088/1742-5468/ab9e64}

\bibitem{hen2018}
I.~Hen,
  \href{https://iopscience.iop.org/article/10.1088/1742-5468/aabbe4}{Off-diagonal
  series expansion for quantum partition functions}, J. Stat. Mech. 2018~(5)
  (2018) 053102.
\newblock \href {https://doi.org/10.1088/1742-5468/aabbe4}
  {\path{doi:10.1088/1742-5468/aabbe4}}.
\newline\urlprefix\url{https://iopscience.iop.org/article/10.1088/1742-5468/aabbe4}

\bibitem{hen2021}
I.~Hen,
  \href{https://link.aps.org/doi/10.1103/PhysRevResearch.3.023080}{Determining
  quantum {M}onte {C}arlo simulability with geometric phases}, Phys. Rev. Res.
  3 (2021) 023080.
\newblock \href {https://doi.org/10.1103/PhysRevResearch.3.023080}
  {\path{doi:10.1103/PhysRevResearch.3.023080}}.
\newline\urlprefix\url{https://link.aps.org/doi/10.1103/PhysRevResearch.3.023080}

\bibitem{akaturk2024}
E.~Akaturk, I.~Hen,
  \href{https://link.aps.org/doi/10.1103/PhysRevB.109.134519}{Quantum {Monte}
  {Carlo} algorithm for {Bose-Hubbard} models on arbitrary graphs}, Phys. Rev.
  B 109 (2024) 134519.
\newblock \href {https://doi.org/10.1103/PhysRevB.109.134519}
  {\path{doi:10.1103/PhysRevB.109.134519}}.
\newline\urlprefix\url{https://link.aps.org/doi/10.1103/PhysRevB.109.134519}

\bibitem{babakhani2025}
A.~Babakhani, L.~Barash, I.~Hen, \href{https://arxiv.org/abs/2503.08039}{A
  quantum {M}onte {C}arlo algorithm for arbitrary high-spin {H}amiltonians},
  preprint arXiv:2503.08039 (2025).
\newline\urlprefix\url{https://arxiv.org/abs/2503.08039}

\bibitem{hen2019}
I.~Hen, \href{https://link.aps.org/doi/10.1103/PhysRevE.99.033306}{Resolution
  of the sign problem for a frustrated triplet of spins}, Physical Review E
  99~(3) (2019) 033306.
\newblock \href {https://doi.org/10.1103/PhysRevE.99.033306}
  {\path{doi:10.1103/PhysRevE.99.033306}}.
\newline\urlprefix\url{https://link.aps.org/doi/10.1103/PhysRevE.99.033306}

\bibitem{chen2021}
Y.-H. Chen, A.~Kalev, I.~Hen,
  \href{https://link.aps.org/doi/10.1103/PRXQuantum.2.030342}{Quantum algorithm
  for time-dependent hamiltonian simulation by permutation expansion}, PRX
  Quantum 2 (2021) 030342.
\newblock \href {https://doi.org/10.1103/PRXQuantum.2.030342}
  {\path{doi:10.1103/PRXQuantum.2.030342}}.
\newline\urlprefix\url{https://link.aps.org/doi/10.1103/PRXQuantum.2.030342}

\bibitem{kalev2021quantum}
A.~Kalev, I.~Hen, \href{https://doi.org/10.22331/q-2021-04-08-426}{Quantum
  algorithm for simulating hamiltonian dynamics with an off-diagonal series
  expansion}, Quantum 5 (2021) 426.
\newblock \href {https://doi.org/10.22331/q-2021-04-08-426}
  {\path{doi:10.22331/q-2021-04-08-426}}.
\newline\urlprefix\url{https://doi.org/10.22331/q-2021-04-08-426}

\bibitem{kalev2024}
A.~Kalev, I.~Hen,
  \href{https://link.aps.org/doi/10.1103/PhysRevResearch.7.013220}{Feynman path
  integrals for discrete-variable systems: Walks on hamiltonian graphs}, Phys.
  Rev. Res. 7 (2025) 013220.
\newblock \href {https://doi.org/10.1103/PhysRevResearch.7.013220}
  {\path{doi:10.1103/PhysRevResearch.7.013220}}.
\newline\urlprefix\url{https://link.aps.org/doi/10.1103/PhysRevResearch.7.013220}

\bibitem{kalev2021dyson}
A.~Kalev, I.~Hen, \href{https://dx.doi.org/10.1088/1367-2630/ac2dae}{An
  integral-free representation of the dyson series using divided differences},
  New Journal of Physics 23~(10) (2021) 103035.
\newblock \href {https://doi.org/10.1088/1367-2630/ac2dae}
  {\path{doi:10.1088/1367-2630/ac2dae}}.
\newline\urlprefix\url{https://dx.doi.org/10.1088/1367-2630/ac2dae}

\bibitem{ezzell2025fidsus}
N.~Ezzell, I.~Hen, \href{https://arxiv.org/abs/2408.03924}{A universal
  black-box quantum monte carlo approach to quantum phase transitions},
  preprint arXiv:2408.03924 (2025).
\newline\urlprefix\url{https://arxiv.org/abs/2408.03924}

\bibitem{ezzell2025advanced}
N.~Ezzell, I.~Hen, \href{https://arxiv.org/abs/2504.07295}{Advanced measurement
  techniques in quantum monte carlo: The permutation matrix representation
  approach}, preprint arXiv:2504.07295 (2025).
\newline\urlprefix\url{https://arxiv.org/abs/2504.07295}

\bibitem{KarlinRinott1980}
S.~Karlin, Y.~Rinott,
  \href{https://doi.org/10.1016/0047-259X(80)90065-2}{Classes of orderings of
  measures and related correlation inequalities. i. multivariate totally
  positive distributions}, Journal of Multivariate Analysis 10~(4) (1980)
  467--498.
\newblock \href {https://doi.org/10.1016/0047-259X(80)90065-2}
  {\path{doi:10.1016/0047-259X(80)90065-2}}.
\newline\urlprefix\url{https://doi.org/10.1016/0047-259X(80)90065-2}

\bibitem{roberts1974}
A.~W. Roberts, D.~E. Varberg, Convex Functions, Academic Press, 1974.

\bibitem{kunz1965inverse}
K.~Kunz, \href{https://doi.org/10.1109/PROC.1965.3931}{Inverse laplace
  transforms in terms of divided differences}, Proceedings of the IEEE 53~(6)
  (1965) 617.
\newblock \href {https://doi.org/10.1109/PROC.1965.3931}
  {\path{doi:10.1109/PROC.1965.3931}}.
\newline\urlprefix\url{https://doi.org/10.1109/PROC.1965.3931}

\end{thebibliography}

\documentclass[preprint,12pt]{elsarticle}
\usepackage[T1]{fontenc}
\usepackage[utf8]{inputenc}
\usepackage[english]{babel}
\usepackage{amsmath,amssymb,amsthm,mathtools}
\usepackage{geometry}
\usepackage{microtype}
\usepackage{enumitem}
\usepackage{hyperref}
\usepackage{cleveref}
\crefname{equation}{eq.}{eqs.}
\Crefname{equation}{Eq.}{Eqs.}
\hypersetup{colorlinks=true,linkcolor=blue,urlcolor=blue,citecolor=blue}
\newtheorem{theorem}{Theorem}
\newtheorem{lemma}{Lemma}
\newtheorem{corollary}{Corollary}

\theoremstyle{remark}

\newcommand{\e}{{e}}
\newcommand{\ba}{\begin{eqnarray}}
\newcommand{\ea}{\end{eqnarray}}
\newcommand{\bea}{\begin{eqnarray}}
\newcommand{\eea}{\end{eqnarray}}
\newcommand{\beq}{\begin{equation}}
\newcommand{\eeq}{\end{equation}}
\newcommand{\dtau}{\,d\tau}
\newcommand{\dz}{\,dz}
\newcommand{\lap}[1]{\mathcal{L}\left\{ #1 \right\}}
\newcommand{\pdv}[1]{{\frac{\partial}{\partial #1}}}
\begin{document}
\begin{frontmatter}
\title{Inequalities, identities, and bounds for \\  divided differences of the exponential function}
\begin{abstract}
\noindent 
Let $\exp[x_0,x_1,\dots,x_n]$ denote the divided difference of the exponential function.
\begin{enumerate}[label=(\roman*)]
\item We prove that exponential divided differences are log-submodular.
\item We establish the four-point inequality
\[
\exp[a,a,b,c]\,\exp[d,d,b,c]+\exp[b,b,a,d]\,\exp[c,c,a,d]-\exp[a,b,c,d]^2 \ge 0,
\]
for all \(a,b,c,d\in\mathbb{R}\).
\item We obtain sharp two-sided bounds for $\exp[x_0,\dots,x_n]$ at fixed mean and variance;
as a consequence, we derive their large-input asymptotics.
\item We present closed-form identities for divided differences of the exponential function, including a convolution identity and summation formulas for repeated arguments. 
\end{enumerate}
\end{abstract}
\author[aff1]{Qiulin Zeng} 
\author[aff2,aff3]{Nicholas Ezzell}
\author[aff2,aff3]{Arman Babakhani}
\author[aff2,aff3]{Itay Hen}
\author[aff3]{Lev Barash}

\affiliation[aff1]{organization={Department of Mathematics, University of Southern California},
            city={Los Angeles},
            postcode={90089}, 
            state={CA},
            country={USA}} 
\affiliation[aff2]{organization={Department of Physics, University of Southern California},
            city={Los Angeles},
            postcode={90089}, 
            state={CA},
            country={USA}}
\affiliation[aff3]{organization={Information Sciences Institute, University of Southern California},
            city={Marina Del Rey},
            postcode={90262}, 
            state={CA},
            country={USA}}
\end{frontmatter}

\section{Introduction}

Divided differences of the exponential function lie at the intersection of analysis, operator theory, probability, and computation.
Analytically, the Hermite--Genocchi simplex formula represents $\exp[x_0,\dots,x_n]$ as an average of $e^x$ over the convex hull 
of the nodes, yielding positivity and a~priori bounds for interpolation remainders and related 
extremal problems~\cite{mccurdy1984,deboor2005}.
In functional calculus for matrices and operators, divided differences serve as the kernels of double—and for higher variations, 
multiple—operator integrals, providing a concise framework for perturbation theory and operator inequalities
\cite{DaleckiiKrein,BirmanSolomyak1968,BhatiaMatrixAnalysis}.
Via the moment-generating function, these structures connect directly to probabilistic tail bounds~\cite{Chernoff1952,Hoeffding1963,Tropp2012}.

For matrix functions $f(A)$, direct evaluators such as scaling--and--squaring with Pad\'e and the Schur--Parlett scheme reduce to an upper (block) triangular form and fill $f(T)$ from the diagonal outward: this requires two--point divided differences, with repeated--argument limits when diagonal blocks contain repeated eigenvalues~\cite{Parlett1998,Higham2008}; for $f(x)=e^x$ this viewpoint includes block--triangular formulas of Van Loan type~\cite{VanLoan1978}.
Sensitivity and conditioning at first order are governed by the Fr\'echet derivative, which in a diagonalizing basis depends entrywise on the same two--point differences~\cite{Higham2008}; multi--point divided differences enter higher Fr\'echet derivatives used in second--order error estimates, Hessian-based optimization, and uncertainty quantification.

In time integration for stiff ODEs, exponential integrators such as exponential time differencing and exponential Runge--Kutta methods 
are built from the $\varphi$--functions, which are repeated--argument divided differences of $e^z$; assembling the stages typically needs only two--point differences~\cite{CoxMatthews2002,KassamTrefethen2005,HochbruckOstermann2010,NiesenWright2012}.
However, when using Newton/Leja interpolation or related multipoint polynomial/rational schemes 
to approximate the action of the matrix exponential or the \(\varphi\)-functions on a vector, 
both the coefficients and the error terms are naturally expressed in terms of higher–order divided differences~\cite{deboor2005,CaliariKandolfOstermannRainer2016LejaBEA,KandolfOstermannRainer2014LejaResidual,Guettel2013RationalKrylovReview}.
Higher--order divided differences also appear in multiple operator integrals and higher Fr\'echet derivatives used in advanced perturbation analyses~\cite{Peller2006MOI,Peller2016MOISurvey}.

The problem of stable and efficient computation of exponential divided differences is as nontrivial as it is an important topic, because exponentials exacerbate conditioning issues in divided differences.
Recent advances~\cite{barash2020} allow incremental updates (add/remove) with controlled cost, well-suited to adaptive or evolving-node scenarios.
Other works~\cite{mccurdy1984,Caliari2007DividedDifferences,Zivcovich2019} focus on structuring the computations to reduce error amplification, or to permit stable use of these divided differences in exponential integrators with less restrictive time-step choices.

In computational physics, exponential divided differences play a central role in the permutation matrix representation 
quantum Monte Carlo (PMR--QMC) framework, where they serve as configuration weights. 
This framework is universal and effective for problems in many-body quantum physics 
and materials simulation, and have been studied and applied during last decade~\cite{albash2017,barash2024,gupta2020,hen2018,hen2021,barash2020,akaturk2024,babakhani2025,hen2019,chen2021,kalev2021quantum,kalev2024,kalev2021dyson,ezzell2025fidsus,ezzell2025advanced}.

The present work identifies several properties of exponential divided differences that arise naturally in applications and, 
to our knowledge, have not previously been documented or proved in the literature. 

The paper is organized as follows. 
Section~\ref{sec:logsubmod} establishes log-submodularity and supermodularity of exponential divided differences.
In Section~\ref{sec:fourpoint} we a prove a four-point inequality and in
Section~\ref{sec:bounds} we derive sharp two-sided bounds at fixed mean and variance and their large-input asymptotics. 
Section~\ref{sec:identities} is dedicated to several closed-form identities, including a convolution identity and repeated-argument summations. We provide some concluding remarks in Section~\ref{sec:conc}.

\section{Log-submodularity of the divided differences of the exponential function}
\label{sec:logsubmod}

\noindent
Fix nodes \(a_1 \le \cdots \le a_n \in \mathbb{R}\) with \(n \geq 0\), and integers \(p,q \ge 1\). Define
\[
K_n(x,y) := \exp[a_1,\dots,a_n,x^{(p)},y^{(q)}],
\]
the divided difference of the exponential function at these \(n+p+q\) nodes, with superscripts indicating multiplicities.

We prove that \(K_n\) is \emph{log-submodular}.
It follows from $K_n(x,y) > 0$ and Lemma~\ref{lem:equivalence} that log-submodularity is equivalent
to being \emph{totally negative of order~2} (TN\(_2\)):
\begin{equation}
\label{eq:TN2}
K_n(x_1,y_1)\,K_n(x_2,y_2)\ \le\ K_n(x_1,y_2)\,K_n(x_2,y_1)\qquad\text{for all }x_1\le x_2,\ y_1\le y_2.
\end{equation}

In particular, with the symmetric substitution \((x_1,y_1)=(x,x)\), \((x_2,y_2)=(y,y)\) (for \(x\le y\)), \Cref{eq:TN2} gives
\[
K_n(x,x) K_n(y,y) \le K_n(x,y) K_n(y,x).
\]

\begin{lemma}
\label{lem:equivalence}
Let $K:\mathbb{R}^2\to(0,\infty)$. For $u=(u_1,u_2),v=(v_1,v_2)\in\mathbb{R}^2$ set
$u\wedge v:=\big(\min\{u_1,v_1\},\,\min\{u_2,v_2\}\big)$,
$u\vee v:=\big(\max\{u_1,v_1\},\,\max\{u_2,v_2\}\big)$.
The following are equivalent:
\begin{enumerate}[label=(\roman*)]
\item (\emph{Log-submodularity})
\[
\log K(u)+\log K(v)\ \ge\ \log K(u\wedge v)+\log K(u\vee v) \qquad\text{for all }\, u,v\in\mathbb{R}^2.
\]
\item (\emph{Totally negative of order $2$}) 
\[
K(x_1,y_1)\,K(x_2,y_2)\ \le\ K(x_1,y_2)\,K(x_2,y_1)\qquad\text{for all }\, x_1\le x_2,\ y_1\le y_2.
\]
\end{enumerate}
\end{lemma}

\begin{proof}
\smallskip
\noindent\emph{(ii) follows from (i).}
Fix $x_1\le x_2$ and $y_1\le y_2$, and set $u:=(x_2,y_1)$, $v:=(x_1,y_2)$.
Then $u\wedge v=(x_1,y_1)$ and $u\vee v=(x_2,y_2)$, so (i) gives
\[
\log K(x_2,y_1)+\log K(x_1,y_2)\ \ge\ \log K(x_1,y_1)+\log K(x_2,y_2),
\]
which exponentiates to (ii).

\smallskip
\noindent\emph{Conversely, (i) follows from (ii).}
Fix arbitrary $u=(u_1,u_2)$ and $v=(v_1,v_2)$, and set $w:=u\wedge v$ and $z:=u\vee v$.
There are two cases.

\emph{(a)} If $u\le v$ or $v\le u$ (coordinatewise), then $\{w,z\}=\{u,v\}$.
Hence $K(w)K(z)=K(u)K(v)$ and the desired inequality in (i) holds with equality after taking logs.

\emph{(b)} Otherwise, one coordinate increases while the other decreases; without loss of generality assume
$u_1\le v_1$ and $u_2\ge v_2$. Then, we have $(w_1,z_2)=(u_1,u_2)=u$ and $(z_1,w_2)=(v_1,v_2)=v$.
Applying (ii) with $x_1=w_1\le z_1=x_2$ and $y_1=w_2\le z_2=y_2$ yields
\[
K(w_1,w_2)\,K(z_1,z_2)\ \le\ K(w_1,z_2)\,K(z_1,w_2)=K(u)\,K(v).
\]
Taking logs gives
\(
\log K(w)+\log K(z)\le \log K(u)+\log K(v),
\)
which is exactly the inequality in (i).
\end{proof}

\begin{lemma}
\label{lem:HG-AD}
Let $h:[0,1]\to[0,\infty)$ be integrable and define
\begin{equation}
\label{eq:K-def}
K(x,y)\;=\;\int_{0}^{1} h(c)\,K_0(c x,c y)\,dc,
\qquad
K_0(x,y)\;:= \exp[x^{(p)},y^{(q)}].
\end{equation}
Then $K$ is log-submodular (TN\(_2\)): for all $x_1\le x_2$ and $y_1\le y_2$,
\begin{equation}\label{eq:TN2_K}
K(x_1,y_1)\,K(x_2,y_2)\ \le\ K(x_1,y_2)\,K(x_2,y_1).
\end{equation}
\end{lemma}

\begin{proof}
We have
\beq
\label{eq:K0HG}
K_0(x,y) = \frac{1}{(p-1)!(q-1)!} \int_0^1 \tau^{p-1} (1-\tau)^{q-1} e^{\tau x + (1-\tau) y} d\tau.
\eeq
Let $S=[0,1]\times[0,1]$ and equip it with the product measure
$d\mu(c,\tau) = h(c)\; \tau^{p-1}\; (1-\tau)^{q-1}\; dc\; d\tau$.
Endow $S$ with a total order $\preceq$ as follows:
for $u_1=(c_1,\tau_1),\,u_2=(c_2,\tau_2)\in S$ set $u_1\preceq u_2$ iff one of the
following conditions holds:
\[
\text{(i) } c_1\tau_1>c_2\tau_2;\qquad
\text{(ii) } c_1\tau_1=c_2\tau_2\text{ and } c_1>c_2;\qquad
\text{(iii) } c_1\tau_1=c_2\tau_2,\ c_1=c_2\text{ and }\tau_1\ge\tau_2.
\]
This is a total order, and whenever $u_1\preceq u_2$ we have $c_1\tau_1\ge c_2\tau_2$.
Write $u\vee v=\max\{u,v\}$ and $u\wedge v=\min\{u,v\}$ with respect to $\preceq$.

Fix $x_1\le x_2$ and $y_1\le y_2$ and define nonnegative functions on $S$:
\[
\alpha(c,\tau)=e^{\,c(\tau x_1+(1-\tau)y_1)},\quad
\beta(c,\tau)=e^{\,c(\tau x_2+(1-\tau)y_2)},
\]
\[
\gamma(c,\tau)=e^{\,c(\tau x_1+(1-\tau)y_2)},\quad
\delta(c,\tau)=e^{\,c(\tau x_2+(1-\tau)y_1)}.
\]
If $u_1\preceq u_2$, then
\[
\log\frac{\gamma(u_2)\,\delta(u_1)}{\alpha(u_1)\,\beta(u_2)}
=(x_2-x_1)\,\big(c_1\tau_1-c_2\tau_2\big)\ \ge\ 0,
\]
because $x_2\ge x_1$ and the terms involving $y_1,y_2$ cancel. Hence the local hypothesis
of the Four Functions inequality on the totally ordered space $S$ holds:
\[
\alpha(u_1)\,\beta(u_2)\ \le\ \gamma(u_2)\,\delta(u_1)\qquad\text{for all }u_1\preceq u_2\in S,
\]
i.e., $\alpha(u)\beta(v)\le\gamma(u\vee v)\delta(u\wedge v)$ for all $u,v\in S$.

By the \emph{integral Four Functions Theorem} (Theorem 2.1 in~\cite{KarlinRinott1980}),
this integrates to
\[
\Big(\int_S\alpha\,d\mu\Big)\Big(\int_S\beta\,d\mu\Big)
\ \le\
\Big(\int_S\gamma\,d\mu\Big)\Big(\int_S\delta\,d\mu\Big).
\]
But $\int_S\alpha\,d\mu / ((p-1)! (q-1)!) = K(x_1,y_1)$, $\int_S\beta\,d\mu / ((p-1)! (q-1)!) = K(x_2,y_2)$ and similarly
for $\gamma,\delta$, which is exactly Eq.~(\ref{eq:TN2_K}).
\end{proof}

\begin{corollary}
\label{cor:K0-TN2}
For all $x_1\le x_2$ and $y_1\le y_2$,
\[
K_0(x_1,y_1)\,K_0(x_2,y_2)\ \le\ K_0(x_1,y_2)\,K_0(x_2,y_1),
\]
i.e.\ $K_0$ is TN$_2$.
\end{corollary}

\begin{proof}
For $\varepsilon\in(0,1)$ define
\[
h_\varepsilon(c):=\frac{1}{\varepsilon}\,\mathbf{1}_{[\,1-\varepsilon,\,1\,]}(c),
\qquad
K_\varepsilon(x,y):=\int_{0}^{1} h_\varepsilon(c)\,K_0(c x, c y)\,dc .
\]
By Lemma~\ref{lem:HG-AD}, each $K_\varepsilon$ is TN$_2$.

We have
\[
\bigl|K_\varepsilon(x,y)-K_0(x,y)\bigr|
= \left|\frac{1}{\varepsilon}\!\int_{1-\varepsilon}^{1}\!\bigl(K_0(cx,cy)-K_0(x,y)\bigr)\,dc\right|
\le \sup_{c\in[1-\varepsilon,1]} \bigl| K_0(cx,cy) - K_0(x,y)\bigr|.
\]
Since $c\mapsto K_0(cx,cy)$ is continuous at $c=1$, $K_\varepsilon(x,y)\to K_0(x,y)$ as $\varepsilon\to0$. 
Passing to the limit in the TN$_2$ inequality for $K_\varepsilon$ yields the claim for $K_0$.
\end{proof}

\begin{theorem}[Log-submodularity]
\label{thm:KnTN2}
For all $n\in\{0,1,2,\dotsc\}$, $x_1\le x_2$ and $y_1\le y_2$,
\[
K_n(x_1,y_1)\,K_n(x_2,y_2)\ \le\ K_n(x_1,y_2)\,K_n(x_2,y_1),
\]
i.e.\ $K_n$ is TN$_2$.
\end{theorem}
\begin{proof}
By the Hermite--Genocchi formula for divided differences, we have
\begin{equation}\label{eq:Kn-mixture}
K_n(x,y)=\int_{0}^{1} g_n(1-c)\,K_0\big(c x,c y\big)\,dc ,
\end{equation}
where
\begin{equation}
\label{eq:gn}
g_n(t)\;:=\;\int\displaylimits_{\substack{\sum_{i=1}^n s_i=t\\ s_i\ge 0}}
\exp\left(\sum_{i=1}^n s_i a_i\right)\,ds_1\dots ds_n\ \ \ge\ 0.
\end{equation}
Therefore, the desired property for $n \geq 1$ follows from Lemma~\ref{lem:HG-AD}.
The property for $n=0$ is proved in Corollary~\ref{cor:K0-TN2}.

\end{proof}

\begin{theorem}[Supermodularity]
Let \(n\ge 0\) and integers \(p,q\ge 1\). Then
\[
K_n(x_1,y_1)+K_n(x_2,y_2)\ \ge\ K_n(x_1,y_2)+K_n(x_2,y_1)
\qquad\text{for all }x_1\le x_2,\ y_1\le y_2.
\]
\end{theorem}
\begin{proof}
Since $K_n(x,y)$ is twice continuously differentiable, supermodularity is equivalent to 
$\partial_{xy}K_n(x,y)\ge 0$ for all $(x,y)\in\mathbb{R}^2$.
Using the differentiation identity for repeated arguments,
\[
\partial_{xy}K_n(x,y)=p\,q\,\exp[a_1,\ldots,a_n,x^{(p+1)},y^{(q+1)}]\;>\;0.
\]
Hence $K_n$ is supermodular.
\end{proof}

\section{A Four-Point Inequality for Divided Differences}
\label{sec:fourpoint}

Let $\exp[x_0,x_1,\dots,x_n]$ denote the divided differences of the exponential function, and
\bea
f(a,b,c,d) &=& \exp[a,a,b,c]\exp[d,d,b,c] + \exp[b,b,a,d]\exp[c,c,a,d] - \exp[a,b,c,d]^2, \\
h(x,y) &=& \exp[x,x,y,y](x-y)^2 = \e^x + \e^y - 2\exp[x,y].
\eea
Equivalently,
\beq
\label{eq:h-phi}
h(x,y)=2\e^{(x+y)/2}\,\phi\!\left(\frac{x-y}{2}\right),\qquad
\phi(u):=\cosh u-\frac{\sinh u}{u}\quad\text{with}\ \ \phi(0):=0.
\eeq

We prove the four-point inequality
\[
\exp[a,a,b,c]\exp[d,d,b,c] + \exp[b,b,a,d]\exp[c,c,a,d] - \exp[a,b,c,d]^2 \geq 0,
\]
which holds for all $a,b,c,d \in\mathbb{R}$.

\begin{lemma}
\label{lemma:h}
\beq
f(a,b,c,d) = \frac{h(a,c)h(b,d)+h(a,b)h(c,d)-h(b,c)h(a,d)}{2(a-b)(a-c)(b-d)(c-d)}.
\label{eq:statement1}
\eeq
\end{lemma}
\begin{proof}
We have
\begin{multline}
f(a,b,c,d) = 
\frac{\e^a + \exp[b,c] - \exp[a,c] - \exp[a,b]}{(a-b)(a-c)}\times\frac{\e^d + \exp[b,c] - \exp[b,d] - \exp[c,d]}{(b-d)(c-d)}
+ \\ +
\frac{\e^b + \exp[a,d] - \exp[a,b] - \exp[b,d]}{(a-b)(b-d)}\times\frac{\e^c + \exp[a,d] - \exp[a,c] - \exp[c,d]}{(a-c)(c-d)}
- \\ -
\frac{\exp[a,c]-\exp[a,d]-\exp[b,c]+\exp[b,d]}{(a-b)(c-d)}\times\frac{\exp[a,b]+\exp[c,d]-\exp[a,d]-\exp[b,c]}{(a-c)(b-d)}.
\end{multline}
It follows that
\begin{multline}
f(a,b,c,d)(a-b)(a-c)(b-d)(c-d) = 
2\exp[a,c]\exp[b,d] - (\e^a+\e^c)\exp[b,d] - (\e^b+\e^d)\exp[a,c]
+ \\ +
2\exp[a,b]\exp[c,d] - (\e^a+\e^b)\exp[c,d] - (\e^c+\e^d)\exp[a,b] + \e^{b+c} + \e^{a+d} - \\ -
\left(
2\exp[a,d]\exp[b,c] - (\e^b+\e^c)\exp[a,d] - (\e^a+\e^d)\exp[b,c]
\right).
\label{eq:statement1a}
\end{multline}
Eq.~(\ref{eq:statement1}) follows from Eq.~(\ref{eq:statement1a}).
\end{proof}

\begin{lemma}[Triangular inequality]
\beq
h(a,c) \geq h(a,b) + h(b,c) \qquad\text{ when }\quad a\leq b\leq c.
\eeq
\end{lemma}
\begin{proof}
Since $h(x,y) = e^x + e^y - 2\exp[x,y]$, we have
\begin{align*}
h(a,c)-h(a,b)-h(b,c)
&=2\big[(\exp[a,b]-\exp[a,c])+(\exp[b,c]-\exp[b,b])\big]\\
&=2\big[(b-c)\exp[a,b,c]+(c-b)\exp[b,b,c]\big]\\
&=2(c-b)\big(\exp[b,b,c]-\exp[a,b,c]\big)\\
&=2(b-a)(c-b)\,\exp[a,b,b,c]\;\ge\;0.
\end{align*}
Here, we used Newton's recursion for divided differences, the positivity of exponential divided differences, and $a\le b\le c$.
\end{proof}

\begin{lemma}
\label{lem:phi}
\beq
\label{ineq:phi}
\phi(x+y)\phi(y+z)\ \ge\ \phi(x)\phi(z) + \phi(y)\phi(x+y+z) \qquad \text{for all}\quad x,y,z\ge 0.
\eeq
\end{lemma}

\begin{proof}
Let $F(y)=\phi(x+y)\phi(y+z)-\phi(y)\phi(x+y+z)$.
We have 
\begin{equation}\label{eq:Fprime-G}
F'(y)=G_s(y+z)-G_s(y),
\end{equation}
where $s=x+2y+z$ and $G_s(t):=\phi'(s-t)\phi(t)+\phi(s-t)\phi'(t)$ for $t\in[0,s]$.

Since $\phi''(u) = (1+2/u^2)\phi(u)$, we have
\[
G_s'(t) = \phi(s-t)\phi(t)\left(\frac{\phi''(t)}{\phi(t)}-\frac{\phi''(s-t)}{\phi(s-t)}\right) 
= 2\;\phi(s-t)\;\phi(t)\left(t^{-2}-(s-t)^{-2}\right).
\]
Therefore $G_s$ is strictly increasing on $[0,s/2]$, strictly decreasing on $[s/2,s]$,
and attains its maximum at $t=s/2$.
Moreover, $G_s$ is symmetric about $s/2$ since $G_s(s-t)=G_s(t)$.

Recall $s=x+2y+z$. Then
\[
\left|\,y-\frac{s}{2}\,\right|=\frac{x+z}{2},\qquad
\left|\,y+z-\frac{s}{2}\,\right|=\frac{|x-z|}{2}\;\le\;\frac{x+z}{2}.
\]
Thus $y+z$ lies no farther from the maximizer $s/2$ than $y$ does.
By the symmetry and unimodality of $G_s$, this implies
\[
G_s(y+z)\;\ge\;G_s(y).
\]
In view of Eq.~(\ref{eq:Fprime-G}), we have $F'(y)\ge 0$ for all $y\ge 0$.
Since $F$ is nondecreasing on $[0,\infty)$ and $F(0)=\phi(x)\phi(z)$, we obtain Eq.~(\ref{ineq:phi}).
\end{proof}

\begin{lemma}
\label{lemma:hpositive}
For any $a<b<c<d$ one has
\[
h(a,c)h(b,d) \ge h(b,c)h(a,d) + h(a,b)h(c,d).
\]
\end{lemma}

\begin{proof}
Let $x=(b-a)/2$, $y=(c-b)/2$, $z=(d-c)/2$. 
It follows from Eq.~(\ref{eq:h-phi}) that
\beq
h(a,c)h(b,d)-h(b,c)h(a,d)-h(a,b)h(c,d) = 4 e^{(a+b+c+d)/2} Q(x,y,z),
\eeq
where
\beq
Q(x,y,z) = \phi(x+y)\phi(y+z) - \phi(y)\phi(x+y+z) - \phi(x)\phi(z).
\eeq
It follows from Lemma~\ref{lem:phi} that 
\[
Q(x,y,z)\ \ge\ 0,
\]
and therefore the lemma follows.
\end{proof}

\begin{theorem}[Four-point inequality]
For all $a,b,c,d \in \mathbb{R}$ one has
\beq
\label{eq:inequality}
\exp[a,a,b,c]\exp[d,d,b,c] + \exp[b,b,a,d]\exp[c,c,a,d] - \exp[a,b,c,d]^2 \geq 0.
\eeq
\end{theorem}
\begin{proof}
$f(a,b,c,d)$ is invariant under the eight-element subgroup $G$ of the symmetric group $S_4$ acting on $\{a,b,c,d\}$:
$G=\{\mathrm{id},\ (ad),\ (bc),\ (ad)(bc),\ (ab)(cd),\ (ac)(bd),\ (abdc),\ (acdb)\}$.
Hence the $24$ permutations of $(a,b,c,d)$ split into three $G$-orbits.
Let $\alpha<\beta<\gamma<\delta$ be the increasing rearrangement of $\{a,b,c,d\}$.
It suffices to verify that 
$f(\alpha,\beta,\gamma,\delta) \ge0$, $f(\alpha,\gamma,\delta,\beta) \ge0$, and $f(\alpha,\beta,\delta,\gamma) \ge0$.
Since $h(x,y)\ge 0$ for all $x,y\in\mathbb{R}$,
the claim follows directly from Lemmas~\ref{lemma:h} and~\ref{lemma:hpositive} in each case.
If some of the points coincide, the result follows by continuity of divided differences.
\end{proof}

\section{A Sharp Sandwich Bound for Exponential Divided Differences at Fixed Mean and Variance}
\label{sec:bounds}

Let $\Delta_n=\{\lambda\in[0,1]^{n+1}:\ \sum_{i=0}^n\lambda_i=1\}$ be the standard simplex,
and let $\gamma(n,z)=\int_0^z t^{n-1}e^{-t}\,dt$ denote the lower incomplete gamma function.
Throughout we assume $n\ge 1$.
Write
\begin{equation}
\label{eq:notations}
\mu:=\frac1{n+1}\sum_{i=0}^n x_i,\qquad
\sigma^2:=\frac1{n+1}\sum_{i=0}^n (x_i-\mu)^2,\qquad
a_0:=\sqrt{n}\,\sigma,\qquad a:=\frac{n+1}{\sqrt{n}}\,\sigma.
\end{equation}
\begin{equation}
L_n(\sigma) := n\,e^{-a_0}\,(-a)^{-n}\,\gamma\big(n,-a\big),\qquad
M_n(\sigma) := n\,e^{a_0}\,a^{-n}\,\gamma\big(n,a\big).
\end{equation}

\begin{theorem}[Sharp sandwich bound]\label{thm:sandwich}
For any $x_0,\dots,x_n\in\mathbb R$ with $n \ge 1$, 
mean $\mu$ and variance $\sigma^2$ as above,
\begin{equation}\label{eq:sandwich}
e^\mu\, L_n(\sigma)\;\le\;n!\,\exp[x_0,\dots,x_n]\;\le\;e^\mu\, M_n(\sigma).
\end{equation}
Moreover, within the class of vectors with mean $\mu$ and variance $\sigma^2$,
both inequalities are sharp: equality on the right holds for the
configuration $(\mu+a_0,\ \mu-a_0/n,\dots,\mu-a_0/n)$
(up to permutation), and equality on the left holds for the configuration
$(\mu-a_0,\ \mu+a_0/n,\dots,\mu+a_0/n)$ (up to permutation).
\end{theorem}

\begin{proof}
Write $x_i=\mu+y_i$ with $\sum_{i=0}^n y_i=0$ and $\sum_{i=0}^n y_i^2=(n+1)\sigma^2$.
Then, $\exp[x_0,\dots,x_n] = e^\mu\exp[y_0,\dots,y_n]$.
By the Hermite--Genocchi formula,
\begin{equation}
\label{eq:HG}
\exp[y_0,\dots,y_n]
=\int_{\Delta_n} \exp\!\Big(\sum_{i=0}^n \lambda_i y_i\Big)\, d\lambda,
\end{equation}
so the map $y:=(y_0,\dots,y_n)\mapsto \Phi(y):=n!\exp[y_0,\dots,y_n]$ 
is symmetric (permutation invariant) and jointly convex (an integral of convex maps).
Hence it is Schur--convex (see, e.g.,~\cite{roberts1974}, p.~258).
Let
\[
\mathcal{S}\;:=\;\Big\{y\in\mathbb{R}^{n+1}:\ \sum_{i=0}^{n} y_i=0,\ \sum_{i=0}^{n} y_i^2=(n+1)\sigma^2\Big\},
\,\,
y^{+}:=\left(a_0,-\frac{a_0}{n},\dots,-\frac{a_0}{n}\right),\,\, y^{-}:=-y^{+}.
\]

Because $\Phi$ is Schur--convex, Lemma~\ref{lem:twolevel} gives
$\Phi(y^{-})\le \Phi(y)\le \Phi(y^{+})$ for all $y\in\mathcal{S}$, with equality precisely at
the two-level vectors (up to permutation). Since
$\gamma(n,a) = a^n \int_0^1 s^{n-1} e^{-as} ds$,
$\exp[a_0,t] = e^{a_0}\int_0^1 e^{t-a_0} ds$, and
$(n-1)!\exp\big[a_0,t,\dots,t\big]={d^{\,n-1}} \exp[a_0,t] / {d t^{\,n-1}}$, where the node $t$ is repeated $n$ times, we obtain
\begin{eqnarray}
\Phi(y^{+}) = n! \exp\left[a_0,-\frac{a_0}{n},\dots,-\frac{a_0}{n}\right] &=& n e^{a_0} a^{-n} \gamma(n,a), \\
\Phi(y^{-}) = n! \exp\left[-a_0,\,\frac{a_0}{n},\,\dots\,,\,\frac{a_0}{n}\right] &=& n\,e^{-a_0}\,(-a)^{-n}\,\gamma(n,-a).
\end{eqnarray}
Multiplying by $e^\mu$ yields Eq.~(\ref{eq:sandwich}) and the equality cases.
\end{proof}

\begin{lemma}\label{lem:twolevel}
$y^{+}$ is the majorization-maximal vector in $\mathcal{S}$ and $y^{-}$ is the
majorization-minimal vector in $\mathcal{S}$. Equivalently, for every $y\in\mathcal{S}$ with
nonincreasing rearrangement $y_{[1]}\ge\cdots\ge y_{[n+1]}$ and $k=1,\dots,n$,
\[
\sum_{i=1}^k y_{[i]}\ \le\ \sum_{i=1}^k y^{+}_{[i]}
\quad\text{and}\quad
\sum_{i=1}^k y^{-}_{[i]}\ \le\ \sum_{i=1}^k y_{[i]},
\]
with equality for all $k$ iff $y$ is a permutation of $y^{+}$ or $y^{-}$, respectively.
\end{lemma}

\begin{proof}
Write the coordinates of $y\in\mathcal{S}$ in nonincreasing order:
$y_{[1]}\ge y_{[2]}\ge\cdots\ge y_{[n+1]}$. Since $\sum_{i=2}^{n+1} y_{[i]}=-y_{[1]}$, the
Cauchy--Schwarz inequality yields
\[
\sum_{i=2}^{n+1} y_{[i]}^2\ \ge\ \frac{\big(\sum_{i=2}^{n+1} y_{[i]}\big)^2}{n}
\;=\;\frac{y_{[1]}^2}{n}.
\]
Hence
\[
(n+1)\sigma^2=\sum_{i=1}^{n+1} y_{[i]}^2
\ \ge\ y_{[1]}^2+\frac{y_{[1]}^2}{n}
\ =\ y_{[1]}^2\Big(1+\frac1n\Big),
\]
so
\begin{equation}\label{eq:top-cap}
y_{[1]}\ \le\ a_0=\sqrt{n}\,\sigma.
\end{equation}
Moreover, equality in Eq.~(\ref{eq:top-cap}) holds if and only if
$y_{[2]}=\cdots=y_{[n+1]}=-a_0/n$ (equality case in Cauchy--Schwarz), i.e., $y$ is a
permutation of $y^{+}$.

Let $S_k(y):=\sum_{i=1}^k y_{[i]}$ be the $k$-th partial sum. Suppose there existed
$y\in\mathcal{S}$ with $S_k(y)\ge S_k(y^{+})$ for all $k$ and strict inequality for some $k$.
Then, in particular, $S_1(y)=y_{[1]}\ge S_1(y^{+})=a_0$, which by Eq.~(\ref{eq:top-cap}) forces
$y_{[1]}=a_0$ and hence (by the equality case) $y$ is a permutation of $y^{+}$; thus all
partial sums coincide and no strict improvement is possible. Therefore $y^{+}$ is
majorization-maximal in $\mathcal{S}$. The argument applied to $-y$ shows $y^{-}$ is
majorization-minimal.
\end{proof}
\begin{lemma}
\label{lem:LnMn-asymp}
\begin{eqnarray}
\label{eq:Ln}
L_n(\sigma) &=& 1 + \frac{\sigma^2}{2n} - \frac{\sigma^3}{3n^{3/2}} + O\left(\frac{1}{n^2}\right) \qquad \text{as}\ n\to\infty,\\
\label{eq:Mn}
M_n(\sigma) &=& 1 + \frac{\sigma^2}{2n} + \frac{\sigma^3}{3n^{3/2}} + O\left(\frac{1}{n^2}\right) \qquad \text{as}\ n\to\infty.
\end{eqnarray}
\end{lemma}
\begin{proof}
For integer $n\ge1$,
\[
\gamma(n,z)=\sum_{m=0}^\infty \frac{(-1)^m}{(n+m)\,m!}\,z^{\,n+m},
\]
hence
\[
M_n(\sigma)=e^{a_0}\sum_{m=0}^\infty (-1)^m\frac{n}{n+m}\frac{a^m}{m!},\qquad
L_n(\sigma)=e^{-a_0}\sum_{m=0}^\infty \frac{n}{n+m}\frac{a^m}{m!}.
\]

Use the expansion
$n/(n+m)=1-m/n+m^2/n^2-m^3/n^3+O(m^4/n^4)$
and the identities
$\sum_{m=0}^\infty m^k (-a)^m/m!=T^k (e^{-a})$ and
$\sum_{m=0}^\infty m^k a^m/m!=T^k (e^{a})$
for $T:=a\cdot d/da$ and $k=0,1,2,3$.

It follows that
\begin{align*}
\sum_{m=0}^\infty (-1)^m\frac{n}{n+m}\frac{a^m}{m!}
&=e^{-a}\!\left[1+\frac{a}{n}+\frac{a^2-a}{n^2}+\frac{a^3-3a^2+a}{n^3}\right]
+O\!\left(\frac{1}{n^2}\right),\\
\sum_{m=0}^\infty \frac{n}{n+m}\frac{a^m}{m!}
&=e^{a}\!\left[1-\frac{a}{n}+\frac{a^2+a}{n^2}-\frac{a^3+3a^2+a}{n^3}\right]
+O\!\left(\frac{1}{n^2}\right),
\end{align*}
Since $a_0-a=-\sigma/\sqrt n$,
\[
e^{\pm(a_0-a)}=1\pm\frac{\sigma}{\sqrt n}+\frac{\sigma^2}{2n}\pm\frac{\sigma^3}{6n^{3/2}}
+O\!\left(\frac{1}{n^2}\right),
\]
and
\[
\frac{a}{n}=\frac{\sigma}{\sqrt n}+O\!\left(\frac{1}{n^{3/2}}\right),\quad
\frac{a^2-a}{n^2}=\frac{\sigma^2}{n}+O\!\left(\frac{1}{n^{3/2}}\right),\quad
\frac{a^3-3a^2+a}{n^3}=\frac{\sigma^3}{n^{3/2}}+O\!\left(\frac{1}{n^2}\right).
\]
Multiplying and keeping terms up to $n^{-3/2}$ gives Eqs.~(\ref{eq:Ln}) and (\ref{eq:Mn}).
\end{proof}

\begin{corollary}[Large-input limit]\label{thm:ddinfty}
Let $(x_i)_{i=0}^\infty$ be a sequence of real numbers and let
\[
\mu_n:=\frac1{n+1}\sum_{i=0}^n x_i,\qquad
\sigma_n^2:=\frac1{n+1}\sum_{i=0}^n (x_i-\mu_n)^2.
\]
Assume $\sup_{n\ge0}\sigma_n^2\le C<\infty$. Then
\[
n!\,\exp[x_0,\dots,x_n]
=\exp\!\left(\mu_n+\frac{\sigma_n^2}{2n}+O\!\big(n^{-3/2}\big)\right)\quad \text{as } n\to\infty.
\]
\end{corollary}

\section{Useful identities}
\label{sec:identities}

\noindent
In this section, we adopt the shorthand notation $e^{t [x_0, \ldots, x_q]} := f[x_0, \ldots, x_q]$ for $f(x) = e^{t x}$
and $x_0,\dots,x_q\in\mathbb C$.
We begin by briefly recalling several well-known results that will be used below.

The divided difference of any holomorphic function $f(x)$ can be defined over the multiset $[x_0, \ldots, x_q]$ 
using a contour integral~\cite{mccurdy1984, deboor2005},
\begin{equation}
    \label{eq:contour-int-dd}
    f[x_0, \ldots, x_q] \equiv \frac{1}{2\pi i} \oint\limits_{\Gamma} \frac{f(x)}{\prod_{i=0}^q(x - x_i)} \,dx,
\end{equation}
for $\Gamma$ a positively oriented contour enclosing all the $x_i$'s.
The divided differences can be shown to satisfy the Leibniz rule~\cite{deboor2005},
\begin{equation}
    \label{eq:leibniz-rule}
    (f \cdot g)[x_0, \ldots, x_q] = \sum_{j=0}^q f[x_0, \ldots, x_j] g[x_j, \ldots, x_q] = \sum_{j=0}^q g[x_0, \ldots, x_j] f[x_j, \ldots, x_q].
\end{equation}
Replacing the variable $x \rightarrow \alpha x$ in~\Cref{eq:contour-int-dd}, we find the rescaling relation,
\begin{equation}
    \label{eq:rescaling-relation}
    \alpha^q e^{t [\alpha x_0, \ldots, \alpha x_q]} = e^{\alpha t [x_0, \ldots, x_q]}.
\end{equation}
Combining~\Cref{eq:rescaling-relation} with Eq. (11) of Ref.~\cite{kunz1965inverse} with $P_m(x) = 1$, we find
\begin{equation}
    \label{eq:laplace-of-dd}
    \lap{\,e^{\alpha t [x_0, \ldots, x_q]}\,}
    \;=\;
    \frac{\alpha^q}{\prod_{j=0}^q \bigl(s-\alpha x_j\bigr)},
\end{equation}
where $\lap{\cdot}$ denote the Laplace transform in $t \rightarrow s$,
and Eq.~(\ref{eq:laplace-of-dd}) holds in the right half-plane $\Re s \;>\; \max_{0\le j\le q} \Re(\alpha x_j)$,
where the Laplace transform converges.
One can alternatively derive~\Cref{eq:laplace-of-dd} by directly performing the integration to the contour 
integral definition in~\Cref{eq:contour-int-dd}, e.g. by Taylor expanding $e^{\alpha t x}$ and re-summing term-by-term, 
legitimate by the uniform convergence of the exponential.

\begin{theorem}[Convolution]
\label{app-lem:convolution}
Let $j\in\{0,\dots,q-1\}$ and $\beta\ge0$. Then
\begin{equation}
\label{app-eq:convolution-theorem}
\int_0^\beta e^{-\tau [x_{j+1}, \ldots, x_q]}  e^{-(\beta-\tau) [x_0, \ldots, x_j]} \dtau = - e^{-\beta [x_0, \ldots, x_q]}.
\end{equation}
\end{theorem}
\begin{proof}
Define 
\[
f(t):=e^{-t [x_{j+1}, \ldots, x_q]},\qquad 
g(t):=e^{-t [x_0, \ldots, x_j]}.
\]
Their convolution is
\[
(f * g)(t)=\int_0^{t} f(\tau)\, g(t-\tau)\, \dtau,
\]
so Eq.~(\ref{app-eq:convolution-theorem}) states that $(f*g)(\beta)=-e^{-\beta [x_0,\ldots,x_q]}$.
By the convolution property of the Laplace transform and~\Cref{eq:laplace-of-dd}, we find
\begin{multline*}
    \lap{(f * g)(t)} = \lap{f(t)} \lap{g(t)} \\
    = \left( \frac{ (-1)^{q-j-1} }{ \prod_{l=j+1}^q (s + x_l) } \right) \left( \frac{ (-1)^{j} }{ \prod_{m=0}^j (s + x_m) }\right) = \frac{ (-1)^{q-1} }{ \prod_{l=0}^q (s + x_l) } = \lap{ - e^{-t [x_0, \ldots, x_q]} }.
\end{multline*}
By the uniqueness theorem for the Laplace transform, $(f*g)(t)=-e^{-t [x_0,\ldots,x_q]}$ for all $t\ge 0$.
Evaluating at $t=\beta$ gives Eq.~(\ref{app-eq:convolution-theorem}).
\end{proof} 

\begin{theorem}[Repeated argument sum]
    \label{prop:repatedarg-sum-simp}
    \begin{equation}
        \sum_{j=0}^q e^{-\tau [x_0, \ldots, x_q, x_j]} = -\tau e^{-\tau [x_0, \ldots, x_q]}
    \end{equation}
\end{theorem}
\begin{proof}
    Denote $f(\tau) \equiv e^{-\tau [x_0, \ldots, x_q]}$ and recall  $\lap{e^{-\tau [x_0, \ldots, x_q]}} = (-1)^q / (\prod_{i=0}^q (s + x_i)).$ By the differentiation property of the Laplace transform and Eq.~(\ref{eq:laplace-of-dd}),
        \begin{equation}
            \lap{-\tau f(\tau)} = \pdv{s}{\lap{f(\tau)}} = \sum_{j=0}^q \frac{(-1)^{q+1}}{(s + x_j) \prod_{i=0}^q (s + x_i)} = \lap{\sum_{j=0}^q e^{-\tau [x_0, \ldots, x_q, x_j]}},
        \end{equation}
    and the result follows by the uniqueness of the Laplace transform.
\end{proof}

\begin{theorem}[Weighted, repeated argument sum]
\label{lemm:weighted-repeatedarg-sum}
    \beq
        \label{eq:weighted-repeatedarg-sum}
        \sum_{j=0}^q x_j e^{-\tau [x_0,\dots,x_q,x_j]} = \left(\tau\frac{\partial}{\partial\tau}-q\right)e^{-\tau [x_0,\dots,x_q]}.
    \eeq
\end{theorem}

\begin{proof}
We prove this by induction.

\emph{Base case.}
When $q=0$, we have 
$x_0 e^{-\tau [x_0,x_0]} = x_0 \partial e^{-\tau x_0} / \partial x_0 = -\tau x_0 e^{-\tau x_0}$, so Eq.~(\ref{eq:weighted-repeatedarg-sum}) holds.

\emph{Induction step.} By Eq.~(\ref{eq:rescaling-relation}), we write the right-hand side as,
\begin{multline}
    (\tau \partial_\tau - (q + 1)) (-\tau)^{q+1} e^{[-\tau x_0, \ldots, -\tau x_{q+1}]} 
    = -(-\tau)^{q+2} \pdv{\tau} e^{[-\tau x_0, \ldots, -\tau x_{q+1}]} 
 \\
    = -(-\tau)^{q+2} \frac{1}{2\pi i} \oint_{\Gamma}  \pdv{\tau}  \frac{e^z}{\prod_{i=0}^{q+1} (z + \tau x_i)} \dz = \sum_{j=0}^{q+1} x_j e^{-\tau [x_0, \ldots, x_{q+1}, x_j]}
\end{multline}

Therefore, Eq.~(\ref{eq:weighted-repeatedarg-sum}) is true for all $q$.  
\end{proof}

\begin{lemma}[Parametric derivative]
    \begin{equation}
    \label{eq:parametric}
        \pdv{\tau} e^{-\tau [x_0, \ldots, x_q]} = - x_0 e^{-\tau [x_0, \ldots, x_q]} - e^{-\tau [x_1, \ldots, x_q]}
\qquad\text{for}\quad q>0.
    \end{equation}
\end{lemma}
\begin{proof}
    First, define $g(x) = -x e^{-\tau x}$ for convenience. By Eq.~\eqref{eq:contour-int-dd}, 
    \begin{align}
        \pdv{\tau} e^{-\tau [x_0, \ldots, x_q]} \equiv \pdv{\tau} \frac{1}{2\pi i} \oint_{\Gamma} \frac{e^{-\tau z}}{\prod_{i=0}^q (z - x_i)} \dz =  \frac{1}{2\pi i} \oint_{\Gamma} \frac{-z e^{-\tau z}}{\prod_{i=0}^q (z - x_i)} \dz \equiv g[x_0, \ldots, x_q].
    \end{align}
    Employing the Leibniz rule, \Cref{eq:leibniz-rule}, we get the desired result. 
\end{proof}

\begin{corollary}
\beq
    \label{eq:computing-weighted-repeatedarg-sum}
    \sum_{j=0}^q x_j e^{-\tau [x_0,\dots,x_q,x_j]} = 
    	\left\{\begin{array}{ll}
            (-x_0\tau - q)e^{-\tau[x_0,\dots,x_q]}-\tau e^{-\tau[x_1,\dots,x_q]}, & \text{\rm for}\quad q > 0,\\
            -\tau x_0 e^{-\tau x_0}, & \text{\rm for}\quad q = 0.
            \end{array}\right.
\eeq
\end{corollary}
\begin{proof}
    This follows from applying Eq.~(\ref{eq:parametric}) to Theorem~\ref{lemm:weighted-repeatedarg-sum}. 
\end{proof}

\begin{theorem}[Weighted, repeated argument, double, less-than sum]
\beq
\sum_{\substack{i,j=0 \\ i \leq j}}^q x_i e^{-\tau [x_0,\dots,x_q,x_i,x_j]} = 
\left(-\frac{\tau^2}{2}\frac{\partial}{\partial\tau}-\sum_{i=0}^q i\cdot \frac{\partial}{\partial x_i}\right) e^{-\tau [x_0,\dots,x_q]}.
\label{eq:c2a}
\eeq
\end{theorem}

\begin{proof}
We prove this by induction.

Base case.
When $q=0$, we have 
$$
x_0 e^{-\tau [x_0,x_0,x_0]} = x_0 \cdot \frac{1}{2}\frac{\partial^2}{\partial x_0^2} e^{-\tau x_0} 
= \frac{\tau^2 x_0}{2} e^{-\tau x_0} = -\frac{\tau^2}{2}\frac{\partial}{\partial\tau} e^{-\tau x_0},
$$ so Eq.(\ref{eq:c2a}) holds.

Induction step. It follows from Newton's recursion for divided differences that
\begin{align}
\sum_{\substack{i,j=0 \\ i \leq j}}^{q+1} x_i e^{-\tau [x_0,\dots,x_{q+1},x_i,x_j]} &=
\frac{1}{x_0-x_1} \sum_{\substack{i,j=0 \\ i \leq j}}^{q+1} x_i \left( e^{-\tau [x_0,x_2,\dots,x_{q+1},x_i,x_j]} - e^{-\tau[x_1,x_2,\dots,x_{q+1},x_i,x_j]}\right) \notag \\ &= 
\frac{1}{x_0-x_1} \Bigg(
\Big(
\big(-\frac{\tau^2}{2}\frac{\partial}{\partial\tau}-\sum_{i=1}^q i\cdot \frac{\partial}{\partial x_{i+1}}\big)e^{-\tau [x_0,x_2,\dots,x_{q+1}]}  \notag \\
&\qquad + 
x_0 e^{-\tau [x_0,x_2,\dots,x_{q+1},x_0,x_1]}+ 
x_1 \sum_{j=1}^{q+1} e^{-\tau [x_0,\dots,x_{q+1},x_j]}
\Big) \notag \\
-\Big(
\big(&-\frac{\tau^2}{2}\frac{\partial}{\partial\tau}-\sum_{i=1}^q i\cdot \frac{\partial}{\partial x_{i+1}}\big)e^{-\tau [x_1,x_2,\dots,x_{q+1}]} + x_0 \sum_{j=0}^{q+1} e^{-\tau [x_0,x_1,\dots, x_{q+1},x_j]}
\Big)
\Bigg) 
\notag \\ &=
\left(-\frac{\tau^2}{2}\frac{\partial}{\partial\tau}-\sum_{i=0}^q i\cdot \frac{\partial}{\partial x_{i+1}}\right)e^{-\tau [x_0,\dots,x_{q+1}]}
-\sum_{j=1}^{q+1} e^{-\tau [x_0,\dots,x_{q+1},x_j]} \notag \\ &=
\left(-\frac{\tau^2}{2}\frac{\partial}{\partial\tau} -\sum_{i=0}^{q+1} i\cdot \frac{\partial}{\partial x_i}\right) e^{-\tau [x_0,\dots,x_{q+1}]}.
\end{align}

Therefore, Eq.~(\ref{eq:c2a}) is true for all $q$.  
\end{proof}

\begin{corollary}
\begin{multline}
\sum_{\substack{i,j=0 \\ i \leq j}}^q x_i e^{-\tau [x_0,\dots,x_q,x_i,x_j]} \\
	= \left\{\begin{array}{ll}
	\left(\tau^2 x_0/2\right) e^{-\tau [x_0,\dots,x_q]} + (\tau^2/2) e^{-\tau [x_1,\dots,x_q]}
	- \sum_{i=1}^q i \cdot e^{-\tau [x_0,\dots,x_q,x_i]}, & \text{\rm for}\quad q > 0,\\
        (\tau^2 x_0/2) e^{-\tau x_0}, & \text{\rm for}\quad q = 0.
        \end{array}\right.
\label{eq:c3a}
\end{multline}
\end{corollary}
\begin{proof}
Eq.~(\ref{eq:c3a}) follows from Eq.~(\ref{eq:parametric}) and Eq.~(\ref{eq:c2a}).
\end{proof}

\section{Concluding remarks\label{sec:conc}}

The results presented here highlight several structural features of exponential divided differences 
that arise in a broad range of analytic and computational settings.

The log-submodularity and four-point inequality established in this work reveal a latent convexity structure underlying the exponential function's divided differences.
The sharp two-sided bounds at fixed mean and variance provide quantitative control and may prove useful for stability and sensitivity analyses of exponential integrators and operator inequalities.
The closed-form convolution and repeated-argument identities extend the analytic toolkit available for both theoretical investigations and numerical algorithms involving the exponential function and its divided differences, the matrix exponential, and operator analogues.

Beyond their intrinsic mathematical interest, these results suggest new connections among interpolation theory, matrix analysis, and probabilistic inequalities through the common language of exponential divided differences.
Future work may explore analogous properties for other operator-convex functions, extend these inequalities to matrix- or operator-valued settings, and examine their implications for higher-order exponential integrators and quantum Monte Carlo methods.

\section*{Acknowledgments}
N.E. acknowledges partial support from the Graduate Student Fellowship awarded by the USC Department of Physics and Astronomy in the Dornsife College of Letters, Arts, and Sciences, the the Gold Family Fellowship, the ARO MURI grant W911NF-22-S-000, and the U.S. Department of Energy (DOE) Computational Science Graduate Fellowship under Award No. DE-SC0020347 during parts of this work.
I.H. acknowledges support by the Office of Advanced Scientific Computing Research of the U.S. Department of Energy under Contract No. DE-SC0024389. 

\bibliographystyle{elsarticle-num}
\bibliography{main-v4-refs}

\end{document}